English

The probability that a machine develops a fault within the first 3 years of use is 0.003. If 40 machines are selected at random,

Advertisements
Advertisements

Question

The probability that a machine develops a fault within the first 3 years of use is 0.003. If 40 machines are selected at random, calculate the probability that 38 or more will not develop any faults within the first 3 years of use.

Sum
Advertisements

Solution

Let X = number of machines that develop a fault.

p = probability that a machine develops a fault within the first 3 years of use

∴ p = 0.003

and q = 1 - p = 1 - 0.003 = 0.997

Given: n = 40

∴ X ~ B (40, 0.003)

The p.m.f. of X is given by

P(X = x) = `"^nC_x  p^x  q^(n - x)`, x = 0, 1, 2,...,n

i.e. p(x) = `"^40C_x  (0.003)^x  (0.997)^(40 - x)`, x = 0, 1, 2, ....,40

P(38 or more machines will develop any fault)

= P(X ≥ 38) = P(X = 38) + P(X = 39) + P(X = 40)

= p(38) + p(39) + p(40)

`= ""^40C_38 (0.003)^38 (0.997)^(40 - 38) + ""^40C_39 (0.003)^39 (0.997)^(40 - 39) + "^40C_40 (0.003)^40 (0.997)^0`

`= (40 xx 39)/(2 xx 1) (0.003)^38 (0.997)^2 + 40(0.003)^39 (0.997)^1 + 1 * (0.003)^40 (0.997)^0`

`= (780)(0.003)^38 (0.997)^2 + (40) (0.003)^39 (0.997) + 1 xx (0.003)^40 xx 1`

`= (0.003)^38 [(780)(0.997)^2 + 40(0.003)(0.997) + (0.003)^2]`

`= (0.003)^38 [775.327 + 0.1196 + 0.000009]`

`= (0.003)^38 (775.446609)`

`= (775.446609)(0.003)^38`

`≈ (775.44)(0.003)^38`

Hence, the probability that 38 or more machines will not develop the fault within 3 years of use = ` (775.44)(0.003)^38`.

shaalaa.com
  Is there an error in this question or solution?
Chapter 8: Binomial Distribution - Miscellaneous exercise 2 [Page 254]

APPEARS IN

Balbharati Mathematics and Statistics 2 (Arts and Science) [English] Standard 12 Maharashtra State Board
Chapter 8 Binomial Distribution
Miscellaneous exercise 2 | Q 12 | Page 254

Video TutorialsVIEW ALL [1]

RELATED QUESTIONS

The probability that a certain kind of component will survive a check test is 0.6. Find the probability that exactly two of the next four components tested will survive.


A die is thrown 6 times. If ‘getting an odd number’ is a success, find the probability of 5 successes. 


A die is thrown 6 times. If ‘getting an odd number’ is a success, find the probability of at most 5 successes.


Five cards are drawn successively with replacement from a well-shuffled deck of 52 cards; find the probability that all the five cards are spades.


Five cards are drawn successively with replacement from a well-shuffled deck of 52 cards, find the probability that only 3 cards are spades


In a box of floppy discs, it is known that 95% will work. A sample of three of the discs is selected at random. Find the probability that none of the floppy disc work.


In a box of floppy discs, it is known that 95% will work. A sample of three of the discs is selected at random. Find the probability that exactly one floppy disc work.


In a box of floppy discs, it is known that 95% will work. A sample of three of the discs is selected at random. Find the probability that exactly two floppy disc work.


Choose the correct option from the given alternatives:

A die is thrown 100 times. If getting an even number is considered a success, then the standard deviation of the number of successes is ______.


Choose the correct option from the given alternatives:

The mean and the variance of a binomial distribution are 4 and 2 respectively. Then the probability of 2 successes is


If X ~ B(4, p) and P(X = 0) = `16/81`, then P(X = 4) = ______.


Choose the correct option from the given alternatives:

The probability of a shooter hitting a target is `3/4` How many minimum numbers of times must he fire so that the probability of hitting the target at least once is more than 0·99?


If the mean and variance of a binomial distribution are 18 and 12 respectively, then n = ______.


Let X ~ B(10, 0.2). Find P(X = 1).


Let X ~ B(10, 0.2). Find P(X ≥ 1).


Let X ~ B(10, 0.2). Find P(X ≤ 8).


The probability that a bomb will hit a target is 0.8. Find the probability that out of 10 bombs dropped, exactly 2 will miss the target.


The probability that a mountain-bike travelling along a certain track will have a tyre burst is 0.05. Find the probability that among 17 riders: exactly one has a burst tyre


The probability that a mountain-bike travelling along a certain track will have a tyre burst is 0.05. Find the probability that among 17 riders: two or more have burst tyre.


The probability that a machine will produce all bolts in a production run within specification is 0.998. A sample of 8 machines is taken at random. Calculate the probability that 7 or 8 machines.


The probability that a machine will produce all bolts in a production run within specification is 0.998. A sample of 8 machines is taken at random. Calculate the probability that at most 6 machines will produce all bolts within specification. 


A computer installation has 10 terminals. Independently, the probability that any one terminal will require attention during a week is 0.1. Find the probabilities that 0.


A computer installation has 10 terminals. Independently, the probability that any one terminal will require attention during a week is 0.1. Find the probabilities that 1.


A computer installation has 10 terminals. Independently, the probability that any one terminal will require attention during a week is 0.1. Find the probabilities that 2.


It is observed that it rains on 12 days out of 30 days. Find the probability that it rains exactly 3 days of week.


If the probability of success in a single trial is 0.01. How many trials are required in order to have a probability greater than 0.5 of getting at least one success?


In Binomial distribution if n is very large and probability success of p is very small such that np = m (constant) then _______ distribution is applied.


Solve the following problem:

An examination consists of 5 multiple choice questions, in each of which the candidate has to decide which one of 4 suggested answers is correct. A completely unprepared student guesses each answer completely randomly. Find the probability that,

  1. the student gets 4 or more correct answers.
  2. the student gets less than 4 correct answers.

In a Binomial distribution with n = 4, if 2P(X = 3) = 3P(X = 2), then value of p is ______.


Choose the correct alternative:

A sequence of dichotomous experiments is called a sequence of Bernoulli trials if it satisfies ______


If X follows a binomial distribution with parameters n = 10 and p. If 4P(X = 6) = P(X = 4), then p = ______ 


In a binomial distribution `B(n, p = 1/4)`, if the probability of at least one success is greater than or equal to `9/10`, then n is greater than ______.


In a binomial distribution `B(n, p = 1/4)`, if the probability of at least one success is greater than or equal to `9/10`, then n is greater than ______.


If X∼B (n, p) with n = 10, p = 0.4 then E(X2) = ______.


In a binomial distribution, n = 4 and 2P(X = 3) = 3P(X = 2), then q = ______.


A pair of dice is thrown 3 times. If getting a doublet is considered a success, find the probability of getting at least two success.

Solution:

A pair of dice is thrown 3 times.

∴ n = 3

Let x = number of success (doublets)

p = probability of success (doublets)

∴  p = `square`, q = `square`

∴ x ∼ B (n, p)

P(x) = nCxpx qn–x

Probability of getting at least two success means x ≥ 2.

∴ P(x ≥ 2) = P(x = 2) + P(x = 3)

= `square` + `square`

= `2/27`


If X is a binomial variable with range {0, 1, 2, 3, 4} and P(X = 3) = 3P(X = 4) then the parameter ‘p’ of the binomial distribution is


Share
Notifications

Englishहिंदीमराठी


      Forgot password?
Use app×